ASAP! What is Ka for the weak acid, HA, if a 0.020 M solution of the...

ASAP!
What is Ka for the weak acid, HA, if a 0.020 M solution of the acid has a pH of 3.29 at 25ºC?

a. 5.1 × 10-2

b. 6.9 × 10-2

c. 2.6 × 10-4

d. 1.3 × 10-5

e. 1.0 × 10-6

What is the conjugate acid of H2PO4(aq)?

a. H3O+

b. H3PO4

c. HPO42–

e. PO43–
Answer #1

th 0 HA (aq) + H20 (1) = Ā (ae) + H2O (aq) I 0.020 0 0 c -a E (0.020-0) Ka = [a) (wzot) min [la] (0.020-n) Given, pH = -log